Quinton is fed up with his supervisor, Wally Binghampton, and decides his health would improve if he retires early. Quinton is 62 years old, and his normal Social Security retirement age is 67. If he retires today, and his normal age retirement PIA is $2,000, how much can Quinton expect to receive as a monthly retirement benefit?

$1,000.

$1,250.

$1,400.

$1,500.
Answer #1

As Quinton is retiring at 62, ie: 60 months before his actual retirement,

his monthly retirement benefit will be = PIA - (5/9 * 1%* 36 Months * 2000) - (5/12 * 1%* 24 Months * 2000)

his monthly retirement benefit will be = 2000 - 400 - 200 = $1400 Option C